Two batteries P and Q each of Emf 8 V and 3 V have internal resistance of 2 Ω and 4 Ω respectively.
Calculate the values of the current through the circuit.
1
Expert's answer
2020-05-29T09:46:37-0400

Assume that the elements are connected in series. So,


total "e.m.f.=8+3=11 V"


total internal resistance "r=2+4=6\Omega"


current "I=11\/6=1.83A".